OpenStudy (anonymous):

find dy/dx of y=x^3lnx

OpenStudy (turingtest):

product rule: y=x^3lnx y'=3x^2lnx+x^3(1/x)=3x^2lnx+x^2=x^2(3lnx+1)

OpenStudy (turingtest):

\[y'=3x^2 \ln x+x^3(1/x)=3x^2 \ln x+x^2=x^2(3\ln x+1)\]